A contribution margin income statement organizes costs by behavior (variable or fixed), rather than by function (operating, selling, or administrative). The contribution margin is the difference between sales and variable expenses.

Byron Manufacturing has one product that sells for $24.00 per unit. The company estimates fixed costs at $6,000, direct materials at $4.00 per unit, direct labor at $5.00 per unit, and variable overhead costs at $3.00 per unit.

Fill in the contribution margin income statement when 730 units are sold:

Byron Manufacturing
Contribution Margin Income Statement
Sales $_______
Less: Variable costs _______
Contribution margin $ ________
Less: Fixed costs _________
Operating income $ ________

Calculate Byron Manufacturing's per unit contribution margin : $_________.

The contribution margin ratio is 50% .

Calculating the Break-even Point:

The break-even point in sales dollars is $___________ which is a break-even point in units of____________ units.

CVP analysis is used to analyze the effects of changes in selling prices, costs and volume on profits. It is also used to determine target profit, the margin of safety, operating leverage, product mix and choosing among marketing strategies and others.

Suppose Byron management has a target operating income of $3,000. Assume the same costs as above and the sell price remains at $24 per unit. How many units does Byron need to sell to meet this goal?

__________ units

What is Byron's margin of safety in sales and in units when Byron sells 730 units?

Margin of safety in sales $__________ Margin of safety in units __________ units

What is the degree of operating leverage when 730 units are sold? If required, round your answer to two decimal places. Answer: _________
Answer #1
Byron Manufacturing
Contribution Margin Income Statement
Sales (730 x $24) 17520
Less: Variable costs [730 x ($4 + $5 + $3)] 8760
Contribution margin 8760
Less: Fixed costs 6000
Operating income 2760

Per unit contribution margin = $24.00 - $12.00 = $12.00

The break-even point in sales dollars is $12000 which is a break-even point in units of 500 units.

Break-even point in sales dollars = Fixed costs/Contribution margin ratio = $6000/50% = $12000

Break-even point in units = $12000/$24 = 500 units

Unit sales required = (Fixed costs + Target operating income)/Contribution margin per unit = ($6000 + $3000)/$12 = $9000/$12 = 750

Margin of safety in sales $5520. Margin of safety in units 230 units.

Margin of safety in sales = Actual sales - Break-even sales = $17520 - $12000 = $5520

Margin of safety in units = 730 - 500 = 230

Degree of operating leverage = Contribution margin/Operating income = $8760/$2760 = 3.17